OpenStudy (abhisar):

A uniform rope of mass 0.1 Kg and length 2.45 m hangs from a ceiling. Find speed of transverse wave in rope at a point 0.5 m distance from the lower end.

Parth (parthkohli):

OK, so I hope you know the formula. That's all that is needed for this.

OpenStudy (abhisar):

\(\sf Velocity=\large \sqrt {\frac{Tension}{mass}}\) ... this one?

Parth (parthkohli):

Yes. To be more precise,\[v = \sqrt\frac{T}{\mu}\]that is \(\mu\) = mass per unit length

OpenStudy (abhisar):

Oh...Yes! One min..

OpenStudy (abhisar):

Thanks .... !!

Parth (parthkohli):

OK, did you calculate tension etc.?

OpenStudy (abhisar):

Yes....Tension will be 1/49 ...answer will be 2.21 m/s

OpenStudy (abhisar):

I was taking the whole mass of rope but as u corrected it should be mass per unit length.

Parth (parthkohli):

\[\sqrt{\frac{1/49}{0.1/2.45}}=2.21~?\]

OpenStudy (abhisar):

Wait..that's not giving 2.21??3 Is it?

Parth (parthkohli):

That is because the tension is not 1/49 N.